OpenStudy (anonymous):

Latonya bought flowers for her mother. She spent $45. Each tulip was $1 and each rose was $2. She bought 3 more tulips than roses. Let t be the number of tulips. Which algebraic sentence would help you figure out how many tulips Latonya bought?

OpenStudy (anonymous):

nice abs

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

not mines

OpenStudy (anonymous):

i no

OpenStudy (anonymous):

um it would be . . .

OpenStudy (anonymous):

lol.

OpenStudy (anonymous):

t=45-2

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

A. t + 2(t + 3) = 45 B. t + 2(t – 3) = 45 C. 2t (t + 3) = 45 D. -3t + 2t = 45

OpenStudy (anonymous):

3t+2t=45

OpenStudy (anonymous):

that is probably going to be the answer

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

you sure

OpenStudy (anonymous):

if i punched his face would randy orton kill me

OpenStudy (anonymous):

idk but he is a wrestler<3

OpenStudy (anonymous):

then yes

OpenStudy (anonymous):

that is (hopefully) the answer

OpenStudy (anonymous):

im takinn a test

OpenStudy (anonymous):

omfg

OpenStudy (anonymous):

we need to work on it more

OpenStudy (anonymous):

k

OpenStudy (anonymous):

well i need to hurry

OpenStudy (anonymous):

hold on go onto the next question

OpenStudy (anonymous):

i cant

OpenStudy (anonymous):

O M G ...

OpenStudy (anonymous):

it is gonna have to be d